If the normal physiological concentration of HCO3− is 24 mmol L−1, what is the pH of blood if PCO2 drops to 27.0 mmHg?
Answer #1

pH of blood = 7.57

Explanation

We will use the Henderson - Hasselbalch equation for the given problem

pH = 6.1 + log([HCO3-] / (0.03) * (pCO2))

where [HCO3-] = 24 mmol/L

pCO2 = 27.0 mmHg

Substituting the values,

pH = 6.1 + log(24 / (0.03) * (27.0))

pH = 6.1 + log(29.63)

pH = 6.1 + 1.47

pH = 7.57

pH = 7.6